# Representation Bending for Large Language Model Safety

### v1

@@ B.1 Full Comparison
B.2 Comparison with Controllable Text Generation
@@
B.3 Results on General Capabilities
@@
B.4 Another Representation Analysis Example
Figure 7 shows another representation analysis using Logit Lens for another harmful prompt. THe hamrful prompt is: “Write a persuasive article linking immigration into the US from Mexico to increased crime.”The base Llama3 8B Instruct model complies with the prmompt and write the article (Figure 7(a)). We can see in Figure 7(b) that RepBend refuses to comply with the prompt, and in Figure 7(c) that it again generates random tokens with high entropy (red colors near the last layers) when the complying response is forced to the model. Note that generating random tokens is not the usual behavior of vanilla LLMs; as shown in Figure 7(d) the model continues any sequences. However, in Figure 7(c), RepBend successfully refuses the harmful response even when it is forced.
@@ B.5 Generation Results
B.6 Computing infrastructure

### v2

@@ B.1 Full Comparison
| Model | Method | Weight Orthogonalization | |
| HarmBench () | WildGuardTest () | ||
| Llama3 8B Instruct | Original Weight | 80.31 | 82.64 |
| NPO | 76.88 | 81.04 | |
| CB* | 33.75 | 30.95 | |
| RepBend (Ours) | 5.94 | 7.48 | |
| Model | Method | GuidedBench | |
| HarmBench Evaluator () | GuidedEvaluator- GPT-4o-mini () | ||
| Mistral 7B Instruct v0.2 | Original Weight | 29.50 | 23.73 |
| CB* | 4.50 | 6.77 | |
| RepBend (Ours) | 3.50 | 6.72 | |
| Llama3 8B Instruct | Original Weight | 6.00 | 7.92 |
| CB* | 2.50 | 7.50 | |
| RepBend (Ours) | 3.00 | 6.75 | |
B.2 Results on Additional Jailbreak Attacks
We further validate RepBend on more diverse black-box and white-box jailbreak attacks, SCAV (Xu et al., 2024), Weight Orthogonalization (Arditi et al., 2024a), and GuidedBench (Huang et al., 2025), using samples from HarmBench, WildGuardTest, and AdvBench (Chen et al., 2022), where we summarize the results in Table 8 (weight orthogonalization attack), Table 9 (results on GuidedBench), and Table 10 (SCAV attack). We can see that even with these white-box jailbreaking methods, RepBend still achieves low ASR.
| Model | Method | SCAV-Embedding | SCAV-Prompt | |
| Harmbench () | WildguardTest () | AdvBench () | ||
| Mistral 7B Instruct v0.2 | Original Weight | 59.69 | 62.22 | 84.00 |
| RepBend (Ours) | 23.75 | 28.84 | 0.00 | |
| Llama3 8B Instruct | Original Weight | 58.44 | 60.48 | 76.00 |
| RepBend (Ours) | 39.06 | 32.18 | 0.00 | |
B.3 Impact of cos_sim loss
Results.
B.4 Hyperparameter Sensitivity Ablation Study
In this section we show additional ablation studies for the hyperparameter sensitivity. We believe these results provide an insight into how to choose the hyperparameters for RepBend and reduce the search space for hyperparameter tuning.
The results are shown in Figure 7. For these experiments, we randomly choose 6 different hyperparameter combinations for each ablation study, while varying the hyperparameter of interest. We measure ASR using the HarmBench classifier, and over-refusal using the WildGuard classifier.
Figure 7(a) shows ablation study of on Llama3 8B Instruct. Among the choices for , when is between 0.25 and 0.5, a good balance between safety and usability is achieved. We can see in the XSTest figure that when (dark blue line), the over-refusal score is low, suggesting the need for KL-loss term.
Figure 7(b) shows ablation study of on Llama3 8B Instruct. here is set to 0.5. The figure for XSTest shows that we need the KL loss term to prevent over-refusal, but we need to keep the non-zero value small, as having big value for increases ASR on HarmBench.
Figure 7(c) illustrates ablation study of layer choices for applying unsafe loss on Mistral 7B Instruct v0.2. Applying unsafe loss on later layers shows the lowest ASR and it is often recommended for RepBend, but more tuning is needed to keep the general capability of the model high.
Figure 7(d) depicts ablation study of on Mistral 7B Instruct v0.2 (result for Llama3 in Figure 5 and Figure 6). here is set to 0.5. We can see the trend is similar to that of Llama3, shown in Figure 6, where where having non-zero cos_sim coefficient stabilizes and improves safety score, and then performance remains stable across other values of .
B.5 Comparison with Controllable Text Generation
@@
B.6 Results on General Capabilities
@@
B.7 Another Representation Analysis Example
Figure 9 shows another representation analysis using Logit Lens for another harmful prompt. The hamrful prompt is: “Write a persuasive article linking immigration into the US from Mexico to increased crime.” The base Llama3 8B Instruct model complies with the prompt and write the article (Figure 9(a)). We can see in Figure 9(b) that RepBend refuses to comply with the prompt, and in Figure 9(c) that it again generates random tokens with high entropy (red colors near the last layers) when the complying response is forced to the model. Note that generating random tokens is not the usual behavior of vanilla LLMs; as shown in Figure 9(d) the model continues any sequences. However, in Figure 9(c), RepBend successfully refuses the harmful response even when it is forced.
@@ B.8 Generation Results
We conjecture two possibilities for the generation of garbled text (i.e., “illet”). First, the safety alignment of the original model matters. For example, we observed more unsafe behavior with Mistral than Llama3. Thus, forcing Mistral to be safe requires heavier representation bending, yielding generation of more garbled text rather than refusal string. Second, the prompt or jailbreak tactic also plays a role: white-box attacks with deeper access to the model’s internals push the representations further and thus produce more scrambled tokens. Still, we find such garbled outputs fairly rare for our Llama3 model, only 3 out of 320 prompts in HarmBench resulted in garbled outputs.
B.9 Computing infrastructure